OpenStudy (anonymous):

How do i solve for x? (1/4)^(x+1) = 8^(2x+3)

sam (.sam.):

\[(1/4)^{x+1} = 8^{2x+3}\] \[4^{-1(x+1)} = 8^{2x+3}\] \[\Large 2^{-2(x+1)} = 2^{3(2x+3)}\] -2(x+1)=3(2x+3) Solve

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

nope, thanks anyways

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

does x = -5/8